3 to the -7th power times 2 to the -4th power
Answer by Boreal(15235)

(1/2187)((1/16)=1/34992=2.8578*10^(-5)
or by logs
-7 log3+-4log2=-4.54397
raise 10 to that power=0.000028578 or 2.8578 x 10^(-5)